40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 49 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  14529   Wed Apr 10 00:33:09 2019 AnjaliUpdateFrequency noise measurementFrequency noise measurement of 1 micron source
  • Attachement #1 shows the input (ch4-green) modulation frequency and the photodiode output (ch1-yellow) when the modulation frequency is about 100 Hz
  • Attachement #2 shows the input (ch4-green) modulation frequency and the photodiode output (ch1-yellow) when the modulation frequency is about 30 Hz
  • The output frequency is varying in accordance with variation in modulation frequency. It is observed that, for a given modulation frequency also, the output frequency is fluctuating. There could be multiple reasons for this behaviour. One of the main reasons is the frequency noise of the laser itself. Also, there could be acoustic noise coupled to the system (eg, by change in length of the fiber).
  • The experimental setup is then modified as shown in attachment #3. The thick beam spliiter is replaced with a thinner one. The mount is also changed such that the transmitted beam can be now coupled to an other photodiode (earlier  the transmitted light was blocked by the mount). One more photodiode (PDA55) is introduced .So now the two photodiodes in the setup are PDA520 and PDA 55. 
  • We then applied frequency modulation on the input laser and observed the output of the two photodiodes. But we didn't get the results as we expected and observed earlier (shown in attachment #1 &2). Looks like, the problem is poor mode matching between the two beams. 
Quote:
  • The alignment of the output beam from the delayed path of MZI to the photodetector was disturbed when we did the polarisation characterisation yesterday. So, today we tried to align the output beam from the delayed path of MZI to the detector .
  • We then observed the beat output from the detector on oscilloscope.We initialy observed a dc shift . We then applied a frequency modulation on the input laser and observed the output on oscilloscope. We expected to see variation in output frequency in accordance with variation of input frequency modulation. But we didnt observe this and we were not really getting the interference pattern. 
  • We tried to make the alignment better. With a better alignment, we could see the interference pattern. We also observed that the output frequency was varying in accordance with variation in the input frequency modulation. We would expect a better result with proper mode matching of the two beams on the photodetector.
Attachment 1: Modulation_frequency_100Hz.jpg
Modulation_frequency_100Hz.jpg
Attachment 2: Modulation_frequency_30Hz.jpg
Modulation_frequency_30Hz.jpg
Attachment 3: Modified_setup.JPG
Modified_setup.JPG
  14528   Tue Apr 9 19:07:12 2019 gautamUpdateALSIR ALS noise budget

Updated the noise budget to include the unsuppressed frequency noise from the EX laser. It does not explain the noise between 10-100 Hz, although the 1-3 Hz noise is close.

Actually, I think the curve that should go on the budget is when the X arm length is locked to the PSL frequency, whereas this is when the X arm is just locally damped. I will update it later tonight.

Update 1010pm: I've uploaded the relevant plot as Attachment #2. Predictably, the unsuppressed frequency noise of the EX laser is now higher, because the MC length is a noisier frequency reference than the arm cavity. But still it is a factor of 10 below the measured ALS noise.

Quote:

Next noises to budget:

  1. In-loop X arm length noise
  2. In-lop EX laser frequency noise
Attachment 1: ALS_noiseBudget.pdf
ALS_noiseBudget.pdf
Attachment 2: ALS_noiseBudget.pdf
ALS_noiseBudget.pdf
  14527   Tue Apr 9 18:44:00 2019 gautamUpdateALSEX Green PDH discriminant measurement

I decided to use the more direct method, of disconnecting feedback to the EX laser PZT, and then looking at the cavity flashes. 

Attachment #1 shows the cavity swinging through two resonances (data collected via oscilloscope). Traces are for the demodulated PDH error signal (top) and the direct photodiode signal (bottom). The traces don't look very clean - I wonder if some saturation / slew rate effects are at play, because we are operating the PD in the 30 dB setting, where the bandwidth of the PD is spec-ed as 260 kHz, whereas the dominant frequency component of the light on the PD is 430 kHz.

The asymmetric horns corresponding to the sideband resonances were also puzzling. Doing the modeling, Attachment #2, I think this is due to the fact that the demodulation phase is poorly set. The PDH modulation frequency is only ~5x the cavity linewidth, so both the real and imaginary parts of the cavity reflectivity contribute to the error signal. If this calculation is correct, we can benefit (i.e. get a larger PDH discriminant) by changing the demod phase by 60 degrees. However, for 230 kHz, it is impractical to do this by just increasing cable length between the function generator and mixer.

Anyway, assuming that we are at the phi=30 degree situation (since the measurement shows all 3 horns going through roughly the same voltage swing), the PDH discriminant is ~40 uV/Hz. In lock, I estimate that there is ~60 uW of light incident on the PDH reflection photodiode. Using the PD response of 0.2 A/W, transimpedance of 47.5 kohm, and mixer conversion loss of 6dB, the shot-noise limited sensitivity is 0.5 mHz/rtHz. The photodiode dark noise contribution is a little lower - estimated to be 0.2 mHz/rtHz. The loop does not have enough gain to reach these levels.

Quote:

PDH discriminant (40 uV/Hz, see this elog) 

Attachment 1: cavityFlashes.pdf
cavityFlashes.pdf
Attachment 2: modelPDH.pdf
modelPDH.pdf
  14526   Tue Apr 9 00:18:19 2019 gautamUpdateALSEX Green PDH error monitor calibrated

wrong assumption - i checked the gain just now, it is G=10, and is running in the "low-noise" mode, so can only drive 4V. fixed elog, filter.

Note: While working at EX, I saw frequent saturations (red led blinking) on the SR560. Looking a the error mon signal on a scope, it had a pk-to-pk amplitude of ~200mV going into the SR560. Assuming the free-swinging cavity length changes by ~1 um at 1 Hz, the green frequency changes by ~15 MHz, which according to the PDH discriminant calibration of 40 uV/Hz should only make a 60 mV pk to pk signal. So perhaps the cavity length is changing by 4x as much, plausible during daytime with me stomping around the chamber I guess.. My point is that if the SR560 get's saturated (i.e. input > 13000 cts), the DQ-ed spectrum isn't trustworthy anymore. Should hook this up to some proper whitening electronics

Quote:

G=10 or G=100?

  14525   Tue Apr 9 00:16:22 2019 ranaUpdateALSEX Green PDH error monitor calibrated

G=10 or G=100?

  14524   Mon Apr 8 23:52:09 2019 gautamUpdateALSEX Green PDH error monitor calibrated

Some time ago, I had done an actuator calibration of ITMX. This suspension hasn't been victim to the recent spate of suspension problems, so I can believe that the results of those measurement are still valid. So I decided to calibrate the in-loop error signal of the EX green PDH loop, which is recorded via an SR560, G=10, by driving a line in ITMY position (thereby modulating the X arm cavity length) while the EX green frequency was locked to the arm cavity length. Knowing the amount I'm modulating the cavity length by (500 cts amplitude sine wave at 33.14159 Hz using awggui, translating to ~17.2 kHz amplitude in green frequency), I demodulated the response in C1:ALS-X_ERR_MON_OUT_DQ channel. At this frequency of ~33 Hz, the servo gain should be large, and so the green laser frequency should track the cavity length nearly perfectly (with transfer function 1/(1+L), where L is the OLG).

The response had amplitude 5.68 +/- 0.10 cts, see Attachment #1. There was a sneaky gain of 0.86 in the filter module, which I saw no reason to keep at this strange value, and so updated to 1, correcting the demodulated response to 6.6 cts. After accounting for this adjustment, the x10 gain of the SR560, and the loop suppression, I put a "cts2Hz" filter in (Attachment #2). I had to guess a value for the OLG at 33 Hz in order to account for the in-loop suppression. So I measured the OLTF using the usual IN1/IN2 method (Attachment #3), and then used a LISO model of the electronics, along with guesses of the cavity pole (18.5 kHz), low-pass filter poles (4x real poles at 70 kHz), PZT actuator gain (1.7 MHz/V) and PDH discriminant (40 uV/Hz, see this elog) to construct a model OLTF. Then I fudged the overall gain to get the model to line up with the measurement between 1-10kHz. Per this model, I should have ~75dB of gain at ~33Hz, so the tracking error to my cavity length modulation should be ~3.05 Hz. Lines up pretty well with the measured value of 4.7 Hz considering the number of guessed parameters. The measured OLG tapers off towards low frequency probably because the increased loop suppression drives one of the measured inputs on the SR785 into the instrument noise floor.

The final calibration number is 7.1 Hz/ct, though the error on this number is large ~30%. Note that these "Hertz" are green frequency changes - so the change to the IR frequency will be half.

Attachment #4 shows the error signal in various conditions, labelled in the legend. Interpretations to follow.

Attachment 1: errMonCalib.pdf
errMonCalib.pdf
Attachment 2: errMon.png
errMon.png
Attachment 3: OLTF.pdf
OLTF.pdf
Attachment 4: EX_frequencyNoises.pdf
EX_frequencyNoises.pdf
  14523   Mon Apr 8 18:28:25 2019 gautamUpdateALSEX Green PDH checkout

I worked on characterizing the green PDH setup at EX, as part of the ALS noise budgeting process. Summary of my findings:

  1. Green doubling efficiency is ~ 1.5 %/W (3mW of green for 450mW of IR). This is ~half of what was measured on the PSL table. There are probably large errors associated with power measurement with the Ophir power meter, but still, seems like a big mismatch.
  2. The green REFL photodiode is a Thorlabs PDA36A
    • It is being run on 30 dB gain setting, corresponding to a transimpedance of 47.5 kohm into high impedance loads. However, the PD bandwidth for this gain setting is 260 kHz according to the manual, whereas the PDH modulation sidebands on the green light are at twice the modulation frequency, i.e. ~560 kHz, so this is not ideal.
    • There was ~250 uW of green light incident on this photodiode, as measured with the Ophir power meter.
    • The DC voltage level was measured to be ~2.7 V on a scope (High-Z), which works out to ~280 uW of power, so the measurements are consistent.
    • When the cavity is locked, there is about 25% of this light incident on the PD, giving a shot noise level of ~25 nV/rtHz. The dark noise level is a little higher, at 40nV/rtHz.
    • Beam centering on the PD looked pretty good to the eye (it is a large-ish active area, ~3mmx3mm).
    • The beam does not look Gaussian at all - there are some kind of fringes visible in the vertical direction in a kind of halo around the main cavity reflection. Not sure what the noise implications of this are. I tried to capture this in a photo, see Attachment #1. Should an Iris/aperture be used to cut out some of this junk light before the reflection photodiode?
  3. The in-going beam was getting clipped on the Faraday Isolator aperture (it was low in pitch).
    • I fixed this by adjusting the upstream steering, and then moving the two PZT mounted green steering mirrors to recover good alignment to the X arm cavity.
    • GTRX level of ~0.5 was recovered.
  4. To estimate the mode-matching of the input beam to the cavity axis, I looked at the reflected light with the cavity locked, and with just the prompt reflection from the ETM:
    • DC light level on the reflection photodiode was monitored using the High-Z input o'scope.
    • Measured numbers are Plocked ~ 660 mV, Pmisaligned ~ 2.6V, giving a ratio of 0.253.
    • While locked, there was a ~ 10 Hz periodic variation in the DC light level on the green REFL photodiode - not sure what was causing this modulation.
    • However, this is inconsistent with a calculation, see Attachment #2. I assumed modulation depth of 90 mrad, round-trip loss of 100 ppm, and Titm = 1.094%, Tetm = 4.579%, numbers I pulled from the core-optics wiki page.
    • Not sure what effect I've missed out on here - to get the model to match the measurement, I have to either assume a higher cavity finesse, or a much higher round-trip loss (5000ppm), both of which seem implausible.

The main motivation was to get the residual frequency noise of the EX laser when locked to the X arm cavity - but I'll need the V/Hz PDH discriminant to convert the in-loop error signal to frequency units. The idea was to look at the PDH error signal on a scope and match up the horn-to-horn voltage with a model to back out said discriminant, but I'll have to double check my model for errors now given the large mismatch I observe in reflected power.

Attachment 1: IMG_7393.JPG
IMG_7393.JPG
Attachment 2: greenModeMatch.pdf
greenModeMatch.pdf
  14522   Mon Apr 8 11:53:17 2019 gautamUpdateCDSc1oaf needs debugging

I tried restarting c1oaf this weekend to see if turning on the MC length FF would affect the ALS noise performance. I burtrestored the filter settings from March 2016. However, I noticed several possible anomalies, which need debugging. I am not turning the model off because of the possibility of having to reboot all the vertex FEs, but this model is totally unusable right now.

  1. Attachment #1 - the vertex seismometer input produces 1e+20 cts at the output of the feedforward filter. Attachment #2 shows the shape of the feedforward filters - doesn't explain the saturation. Since this is a feedforward loop, a runaway loop can't be the explanation either.
  2. The MC length feedforward control signal is supposed to only go to MC2 - but MC1 and MC3 coil outputs were saturated when I enabled the feedforward.
Attachment 1: c1oaf_sat.png
c1oaf_sat.png
Attachment 2: MCL_FF_TF.pdf
MCL_FF_TF.pdf
  14521   Mon Apr 8 00:04:08 2019 gautamUpdateALSIR ALS noise budget

To start the noise budgeting, I decided to measure the "DFD noise", which is really the quadrature sum of the following terms:

  • ZHL-3A (RF amplifier) noise, NF ~ 6dB per spec (~ 1nV/rtHz)
  • Delay line demod board noise, ~30nV/rtHz [measurement]
  • AA board noise [measurement]
  • ADC noise

According to past characterizations of these noises, the ADC noise level, which is expected to be at the level of a few uV/rtHz, is expected to be the dominant noise source.

The measurement was made by disconnecting the NF 1611 free space photodiode from the input to the RF amplifier on the PSL table, and connecting a Marconi (f_carrier = 40 MHz, signal level=-5dBm) instead. The phase tracked output was then monitored, and the resulting digital spectrum is the red curve in Attachment #1. The blue curve is the ASD of fluctuations of the beatnote between the PSL and EX IR beams, as monitored by the DFD system, with the X arm cavity length locked to the PSL frequency via the LSC servo, and the EX green frequency locked to the X arm cavity length by the analog PDH servo. 

Conclusions:

Assuming the Marconi frquency noise is lower than the ones being budgeted:

  1. the measured frequency noise is above the DFD noise - this needs to be budgeted.
  2. The DFD noise level is consistent with a frequency discriminant of 15 uV/rtHz and an ADC noise level of 3 uV/rtHz at high frequencies.

Next noises to budget:

  1. In-loop X arm length noise
  2. In-lop EX laser frequency noise
Attachment 1: DFDnoise.pdf
DFDnoise.pdf
  14520   Sat Apr 6 02:07:40 2019 AnjaliUpdateFrequency noise measurementFrequency noise measurement of 1 micron source
  • The alignment of the output beam from the delayed path of MZI to the photodetector was disturbed when we did the polarisation characterisation yesterday. So, today we tried to align the output beam from the delayed path of MZI to the detector .
  • We then observed the beat output from the detector on oscilloscope.We initialy observed a dc shift . We then applied a frequency modulation on the input laser and observed the output on oscilloscope. We expected to see variation in output frequency in accordance with variation of input frequency modulation. But we didnt observe this and we were not really getting the interference pattern. 
  • We tried to make the alignment better. With a better alignment, we could see the interference pattern. We also observed that the output frequency was varying in accordance with variation in the input frequency modulation. We would expect a better result with proper mode matching of the two beams on the photodetector.
  14519   Fri Apr 5 11:49:30 2019 gautamUpdateALSPSL + X green beat recovery

Since we haven't been using it, the PID control was not enabled on the doubling oven on the PSL table (it is disabled after every power outage event in the lab). I re-enabled it just now. The setpoint according to the label on the TC200 controller is 36.9 C. The PID paramaters were P=250, I=200, D=40. These are not very good as the overshoot when I turned the control on was 44 C, seems too large. The settling time is also too long, after 10 minutes, the crystal temperature as reported by the TC200 front panel is still oscillating. I can't find anything in the elog about what the nominal PID parameter values were. The X end PID seems much better behaved so I decided to try the same PID gains as is implemented there, P=250, I=60, D=25.

With the Ophir power meter, I measured 60mW of IR light going into the doubling oven, 110uW green light coming out, for a conversion efficiency of 2.7%/W, seems pretty great.

Next, I went to EX and tweaked the steering mirror alignment - I wasn't able to improve the transmission significantly using the PZT sliders on the EPICS screen, and the dither alignment servo isn't working. It required quite a substantial common mode yaw shift of the PZT mirrors to make GTRX ~ 0.5. 

Quote:

I plan to recover the green beat note as well and digitize it using the second available DFD channel (eventually for the Y arm) - then we can simultaneously compare the the green and IR performance (though they will have different noise floors as there is less green light on the green beat PDs, and I think lower transimpedance too).

  14518   Fri Apr 5 11:40:57 2019 AnjaliUpdateFrequency noise measurementFrequency noise measurement of 1 micron source
  • Attachment #1 shows the present experimental setup. The photodiode is now replaced with PDA255. The farther end of the fiber (output of the delayed arm) is coupled through a collimator and aligned such that the beam from the delayed path fall on the detector along with the undelayed path of MZI. We tried to measure the frequency noise of the laser with this setup, but we didn’t get anything sensible.
  • One of the main draw backs of the measurement was the polarisation was not aligned properly in the setup. So, then the next step was to identify the polarisation at different locations in the beam path and to maximise the polarisation to either S or P component.

  • So, we introduced HWP at the input beam path after isolator as shown in attachment #1. Also, the polarisation was tested at positions P1, P2, P3, and P4 shown in attachment #1 by placing a polarisation beam splitter at these locations and then by observing the transmitted (P component) and reflected light (S component) using power meter.

  • The observations at different locations are as the follows

Position Input power (mW) P component (mW) S component (mW)
P1 279 145 123
P2 255 113 137
P3 129 67 58
P4 124 66 53

 

  • These observations show that the P and S components are almost equal, and this is not a good polarisation arrangement. At this point, we also had to check whether the incoming beam is linearly polarised or not.

  • To test the same, the PBS was placed at position P1 and the P and S components were observed with power meter as the HWP is rotated.Attachment # 2 shows the results of the same, that is the variation in P and S component as the HWP is rotated.

  • This result clearly shows that the input beam is linearly polarised. The HWP was then adjusted such that the P component is maximum and coupled to the MZI. With this orientation of HWP, the polarisation observed at different positions P1, P2, P3, and P4 are as follows.

Position Input (mW) P component (mW) S component (mW)
P1 283 276 5
P2 248 228 7
P3 126 121 2
P4 128 117 1
  • This shows that the polarisation is linearly polarised as well as it is oriented along the P direction (parallel to the optical table).

  • We have the polarisation maintaining fiber (PM 980) as the delay fiber. The polarisation of the light as it propagates through a PM fiber depends on how well the input beam is coupled to the axis (slow or fast) of the fiber. So, the next task was to couple the light to one of the axes of the fiber.

  • The alignment key on the fiber is a good indication of the axis of the fiber. In our case, the alignment key lines up with the slow axis of the fiber. We decided to couple the light to the fast axis of the fiber. Since the incoming beam is P polarised, the output fiber coupler was  aligned such that the fast axis is parallel to optical table as possible.

  • A PBS was then introduced after the fiber output collimator . There is a HWP (marked as HWP2 in attachment 1) in front of the input coupler of the fiber as well. This HWP was then rotated and observed the P and S component from the PBS that is now placed after the output coupler with a power meter.The idea was , when the light is coupled to the fast axis of the fiber, we will see the maximum at the P componet at the output

  • Attachment # 3 shows the observation. 

  • In this way I tried to find the orientation of the HWP2 such that the P component is maximum at the output. But I was not succeeded in this method and observed that the output was fluctuating when the fiber was disturbed. One  doubt we had was whether the fiber is PM or not . Thus we checked the fiber end with fiber microscope and confirmed that it is PM fiber. 

  • Thus, we modifed the setup as shown in attachement # 4.The photodetector (PDA55) was monitoring the S component and the output of the detector was observed on an oscilloscope. We rotated the HWP2 such that the S component is almost minimum. At the same time, we were disturbing the fiber and was observing whether the output is fluctuating. The HWP2 angle was tweaked around the minimum of S component and observed the output with disturbing the fiber. This way we found the orientation of HWP2  such that the light is coupled to the fast axis of the fiber and the output was not fluctuating while we disturb the fiber. We tested it  by heating the fiber with a heat gun as well and confirmed that the output is not fluctuating and thus the light is coupled to the fast axis of the fiber.

Attachment 1: Modified_experimental_setup.JPG
Modified_experimental_setup.JPG
Attachment 2: Checking_polarisation.pdf
Checking_polarisation.pdf
Attachment 3: Checking_the_polarisation_alignment_of_the_delay_fiber.pdf
Checking_the_polarisation_alignment_of_the_delay_fiber.pdf
Attachment 4: Setup_to_test_the_polarisation_alignment_of_delay_fiber.JPG
Setup_to_test_the_polarisation_alignment_of_delay_fiber.JPG
  14517   Fri Apr 5 01:10:18 2019 gautamUpdateVACTP3 forepump is also noisy

Is this one close to failure as well?

  14516   Fri Apr 5 00:33:58 2019 gautamUpdateALSPromising IR ALS noise

Summary:

I set up a free-space beat on theNW side of the PSL table between the IR beam from the PSL and from EX, the latter brought to the PSL table via ~40m fiber. Initial measurements suggest very good performance, although further tests are required to be sure. Specifically, the noise below 10 Hz seems much improved.

Details:

Attachment #1 shows the optical setup. 

  • I used two identical Thorlabs F220APC collimators to couple the light back into free space, reasoning that the mode-matching would be easiest this way.
  • Only 1 spare K6Xs collimator mount was available (this has the locking nut on the rotational DoF), so I used a K6X for the other mount. The fast axis of the Panda fibers were aligned as best as possible to p-polarization on the table by using the fact that the connector key is aligned to the slow axis.
  • I cut the power coupled into the PSL fiber from ~2.6mW to ~880uW (using a HWP + PBS combo before the input coupling to the fiber) to match the power from EX.
  • The expected signal level from these powers and the NF1611 transimpedance of 700 V/A is ~320 mVpp. After alignment tweaking, I measured ~310mVpp (~ -5dBm) into a 50 ohm input on a scope, so the mode-matching which means the polarization matching and mode overlap between the PSL and EX beams are nearly optimal.
  • To pipe the signal to the delay line electronics, I decided to use the ZHL-3A (G=27dB, 1dB compression at 29.5dBm per spec), so the signal level at the DFD rack was expected (and confirmed via 50 ohm input on o'scope) to be ~19dBm.
  • This is a lot of signal - after the insertion loss of the power splitter, there would still be ~15dBm of signal going to the (nominally 10dBm) LO input of the demod board. This path has a Teledyne AP1053 at the input, which has 10dB gain and 1dBm compression at 26dBm per spec. To give a bit of headroom, I opted on the hacky solution of inserting an attenuator (5dB) in this path - a better solution needs to be implemented.
  • The differential outputs of the demod board go to the CDS system via an AA board (there is no analog whitening).

Yehonathan came by today so I had to re-align the arms and recover POX/POY locking. This alllowed me to lock the X arm length to the PSL frequency, and lock the EX green laser to the X arm length. GTRX was ~0.36, whereas I know it can be as high as 0.5, so there is definitely room to improve the EX frequency noise suppression.

Attachment #2 shows the ALS out-of-loop noise for the PSL+X combo. The main improvements compared to this time last year are electronic. 

  • The failed experiment of making custom I/F amplifier was abandoned and Rich Abbott's original design was reverted to. 
  • New power splitter was installed with 3dB less insertion loss.
  • According to the RF path level monitor, the signal level at the RF input to the demod board is ~10dBm. Per my earlier characterization, this will give us the pretty beefy frequency discriminant of ~15uV/Hz.
  • I estimate the frequency noise of the detection electronics + ADC noise now translate to 1/3 the frequency noise compared to the old system. With some analog whitening, this can be made even better, the electronics noise of the DFD electronics (~50nV/rtHz) is estimated to be <10mHz/rtHz equivalent frequency noise. 
  • Note that the calibration from phase-tracker-servo to units of Hz (~14 kHz / degree) was not changed in the digital system - this should only be a property of the delay line length, and hence, should not have changed as a result of the various electronics changes to the demod board and other electronics.

Next steps:

  • Improve pointing of green beam into X arm cavity.
  • I plan to recover the green beat note as well and digitize it using the second available DFD channel (eventually for the Y arm) - then we can simultaneously compare the the green and IR performance (though they will have different noise floors as there is less green light on the green beat PDs, and I think lower transimpedance too).
Quote:

Mix the beams in free space. We have the beam coming from EX to the PSL table, so once we mix the two beams, we can use either a fiber or free-space PD to read out the beatnote. 

  • This approach means we lose some of the advantages of the fiber based setup (e.g. frequent alignment of the free-space MM of the two interfering beams may be required).
  • Potentially increases sensitivity to jitter noise at the free-space/fiber coupling points
Attachment 1: IMG_7388.JPG
IMG_7388.JPG
Attachment 2: freeSpace_IR_beat.pdf
freeSpace_IR_beat.pdf
  14515   Wed Apr 3 18:35:54 2019 gautamUpdateVACPSL shutter re-opened

PSL shutter was re-opened at 6pm local time. IMC was locked. As of 10pm, the main volume pressure is already back down to the 8e-6 level.

  14514   Wed Apr 3 16:17:17 2019 JonUpdateVACTP2 forepump replaced

I can't explain the mechanical switching sound Gautam reported. The relay controlling power to the TP2 forepump is housed in the main AC relay box under the arm tube, not in the Acromag chassis, so it can't be from that. I've cycled through the pumpdown sequence several times and can't reproduce the effect. The Acromag switches for TP2 still work fine.

In any case, I've made modifications to the vacuum interlocks that will help with two of the issues:

  1. For the "AC power loss" over-triggering: New logic added requiring the UPS to be out of the "on line power, battery OK" state for ~5 seconds before tripping the interlock. This will prevent electrical transients from triggering an emergency shutdown, as seems to be the case here (the UPS briefly isolates the load to battery during such events).
  2. PSL interlocking: New logic added which directly sets C1:AUX-PSL_ShutterRqst --> 0 (closes the PSL shutter) when the main volume pressure is 3 mtorr-500 torr. Previously there was a channel exposed for this interlock (C1:Vac-interlock_high_voltage), but c1aux was not actually monitoring it. Following the convention of every vac interlock, after the PSL shutter has been closed, it has to be manually reopened. Once the pressure is out of this range, the vac system will stop blocking the shutter from reopening, but it will not perform the reopen action itself. gautam: a separate interlock logic needs to be implemented on c1aux (the shutter machine) that only permits the shutter to be opened if the Vac pressure range is okay. The SUS watchdog style AND logic in the EPICS database file should work just fine.

After finishing this vac work, I began a new pumpdown at ~4:30pm. The pressure fell quickly and has already reached ~1e-5 torr. TP2 current and temp look fine.

Quote:

However, when opening V4 (the foreline of TP1 pumped by TP2), I heard a loud repeated click track (~5Hz) from the electronics rack. Shortly after, the interlocks shut down all the TPs again, citing "AC power loss". Something is not right, I leave it to Jon and Chub to investigate.

Attachment 1: IMG_3180.jpg
IMG_3180.jpg
  14513   Wed Apr 3 12:32:33 2019 KojiUpdateALSNote about new fiber couplers

Andrew seems to have an integrated solution of PBS+HWP in a singe mount. Or, I wonder if we should use HWP/QWP before the coupler. I am interested in a general solution for this problem in my OMC setup too.

  14512   Wed Apr 3 10:42:36 2019 gautamUpdateVACTP2 forepump replaced

Bob and Chub concluded that the drypump that serves as TP2's forepump had failed. Steve had told me the whereabouts of a spare Agilent IDP-7. This was meant to be a replacement for the TP3 foreline pump when it failed, but we decided to swap it in while diagnosing the failed drypump (which had 2182 hours continuous running according to the hour counter). Sure enough, the spare pump spun up and the TP2fl pressure dropped at a rate consistent with what is expected. I was then able to spin up TP1, TP2 and TP3. 

However, when opening V4 (the foreline of TP1 pumped by TP2), I heard a loud repeated click track (~5Hz) from the electronics rack. Shortly after, the interlocks shut down all the TPs again, citing "AC power loss". Something is not right, I leave it to Jon and Chub to investigate.

  14511   Wed Apr 3 09:07:46 2019 gautamUpdateVACVac failure

Overnight pressure trends don't suggest anything went awry after the initial interlock trip. Some watchdog script that monitors vacuum pressure and closes the PSL shutter in the event of pressure exceeding some threshold needs to be implemented. Another pending task is to make sure that backup disk for c1vac actually is bootable and is a plug-and-play replacement.

Attachment 1: vacFailOvernight.png
vacFailOvernight.png
  14510   Wed Apr 3 09:04:01 2019 gautamUpdateALSNote about new fiber couplers

The new fiber beam splitters we are ordering, PFC-64-2-50-L-P-7-2-FB-0.3W, have the slow axis working and fast axis blocked. The way the light is coupled into the fibers right now is done to maximize the amount of light into the fast axis. So we will have to do a 90deg rotation if we use that part. Probably the easiest thing to do is to put a HWP immediately before the free-space-to-fiber collimator.

Update 6pm: They have an "SB" version of the part with the slow axis blocked and fast axis enabled, same price, so I'll ask Chub to get it.

  14509   Tue Apr 2 18:40:01 2019 gautamUpdateVACVac failure

While glancing at my Vacuum striptool, I noticed that the IFO pressure is 2e-4 torr. There was an "AC power loss" reported by C1Vac about 4 hours (14:07 local time) ago. We are investigating. I closed the PSL shutter.


Jon and I investigated at the vacuum rack. The UPS was reporting a normal status ("On Line"). Everything looked normal so we attempted to bring the system back to the nominal state. But TP2 drypump was making a loud rattling noise, and the TP2 foreline pressure was not coming down at a normal rate. We wonder if the TP2 drypump has somehow been damaged - we leave it for Chub to investigate and give a more professional assessment of the situation and what the appropriate course of action is.

The PSL shutter will remain closed overning, and the main volume and annuli are valved off. We spun up TP1 and TP3 and decided to leave them on (but they have negligible load).

Attachment 1: vacFail.png
vacFail.png
  14508   Tue Apr 2 15:02:53 2019 JonUpdateCDSITMY freed

blushI renamed all channels on c1susaux2 from "C1:SUS-..." to "C1:SUS2-..." to avoid contention. When the new system is ready to install, those channel names can be reverted with a quick search-and-replace edit.

Quote:

While doing this work, I noticed several errors corresponding to EPICS channel conflicts. Turns out the c1susaux2 EPICS server was left running, and the MEDM screens (and possibly several scripts) were confused. There has to be some other way of testing the new crate, on an isolated network or something - please do not leave the modbus service running as it potentially interferes with normal IFO operation. For good measure, I stopped the process and shut down the machine since I saw nothing in the elog about any running tests.

  14507   Tue Apr 2 14:53:57 2019 gautamUpdateCDSc1vac added to burt

I deleted references to c1vac1 and c1vac2 (which no longer exist) and added c1vac to the autoburt request file list at /opt/rtcds/caltech/c1/burt/autoburt/requestfilelist

  14506   Mon Apr 1 22:33:00 2019 gautamUpdateCDSITMY freed

While Anjali is working on the 1um MZ setup, the pesky ITMY was liberated from the OSEMs. The "algorithm" :

  • Apply a large (-30000 cts) offset to the side coil using the fast system.
  • Approach the zero of the YAW DoF from -2.00V, PIT from +10V (you'll have to jiggle the offsets until the optic is free swinging, and then step the bias down by 0.1). At this point I had the damping off.
  • Once the PIT bias slider reaches -4V, I engaged all damping loops, and brought the optic to its nominal bias position under damping. 

While doing this work, I noticed several errors corresponding to EPICS channel conflicts. Turns out the c1susaux2 EPICS server was left running, and the MEDM screens (and possibly several scripts) were confused. There has to be some other way of testing the new crate, on an isolated network or something - please do not leave the modbus service running as it potentially interferes with normal IFO operation. For good measure, I stopped the process and shut down the machine since I saw nothing in the elog about any running tests.

Quote:

ITMY became stuck during this process

  14505   Mon Apr 1 12:01:52 2019 JonUpdateCDS 

I brought c1susaux back online this morning for suspension-channel test scripting. It had been dead for some time. I followed the procedure outlined in #12542. ITMY became stuck during this process, which Gautam tells me always happens since the last vacuum access, but ITMX is not stuck.

  14504   Sun Mar 31 18:39:45 2019 AnjaliUpdateAUXAUX laser fiber moved from AS table to PSL table
  • Attachment #1 shows the schematic of the experimental setup for the frequency noise measurement of 1 um laser source.

  • AUX laser will be used as the seed source and it is already coupled to a 60 m fiber (PM980). The other end of the fiber was at the AS table and we have now removed it and placed in the PSL table.

  • Attachment # 2 shows the photograph of the experimental setup. The orange line shows the beam that is coupled to the delayed arm of MZI and the red dotted line shows the undelayed path.

  • As mentioned, AUX is already coupled to the 60 m fiber and the other end of the fiber is now moved to the PSL table. This end needs to be collimated. We are planning to take the same collimator from AS table where it was coupled into before. The position where the collimator to be installed is shown in attachment #2. Also, we need to rotate the mirror (as indicated in attachment #2) to get the delayed beam along with the undelayed beam and then to combine them. As indicated in attachment #2, we can install one more photo diode to perform  balanced detection.

  • We need to decide on which photodetector to be used. It could be NF1801 or PDA255.

  • We also performed the power measurement at different locations in the beam path. The different locations at which power measurement is done is shown attachment #3

  • There is an AOM in the beam path that coupled to the delayed arm of MZI. The output beam after AOM was coupled to the zero-order port during this measurement. That is the input voltage to the AOM was at 0 V, which essentially says that the beam after the AOM is not deflected and it is coupled to the zero-order port. The power levels measured at different locations in this condition are as follows. A)282 mW B)276 mW C)274 mW D)274 mW E)273 mW F)278 mW G)278 mW H)261 mW I)263 mW J)260 mW K)131 mW L)128 mW M)127 mW N)130 mW

  • It can be seen that the power is halved from J to K. This because of a neutral density filter in the path of the beam

  • In this case, we measured a power of 55 mW at the output of the delayed fiber. We then adjusted the input voltage to the AOM driver to 1 V such that the output of AOM is coupled to the first order port. This reduced the power level in the zero-order port of AOM that is coupled to the delayed arm of the MZI. In this case we measured a power of 0.8 mW at the output of delayed fiber.

  •  We must be careful about the power level that is reaching the photodetector such that it should not exceed the damage threshold of the detector.

  • The power measured at the output of undelayed path is 0.8 mW.

  • We also must place the QWP and HWP in the beam path to align the polarisation.

Quote:

[anjali, gautam]

To facilitate the 1um MZ frequency stabilization project, I decided that the AUX laser was a better candidate than any of the other 3 active NPROs in the lab as (i) it is already coupled into a ~60m long fiber, (ii) the PSL table has the most room available to set up the readout optics for the delayed/non-delayed beams and (iii) this way I can keep working on the IR ALS system in parallel. So we moved the end of the fiber from the AS table to the SE corner of the PSL table. None of the optics mode-matching the AUX beam to the interferometer were touched, and we do not anticipate disturbing the input coupling into the fiber either, so it should be possible to recover the AUX beam injection into the IFO relatively easily.

Anjali is going to post detailed photos, beam layout, and her proposed layout/MM solutions later today. The plan is to use free space components for everything except the fiber delay line, as we have these available readily. It is not necessarily the most low-noise option, but for a first pass, maybe this is sufficient and we can start building up a noise budget and identify possible improvements.

The AUX laser remians in STANDBY mode for now. HEPA was turned up while working at the PSL table, and remains on high while Anjali works on the layout.

 

Attachment 1: Schematic_of_experimental_setup_for_frequency_stabilisation_of_1_micron_source.png
Schematic_of_experimental_setup_for_frequency_stabilisation_of_1_micron_source.png
Attachment 2: 1_micron_setup_for_frequency_noise_measurement.JPG
1_micron_setup_for_frequency_noise_measurement.JPG
Attachment 3: 1_micron_setup_for_frequency_noise_measurement_power_levels.png
1_micron_setup_for_frequency_noise_measurement_power_levels.png
  14503   Sun Mar 31 15:05:53 2019 gautamUpdateALSFiber beam-splitters not PM

I looked into this a little more today.

  1. Looking at the beat signal between the PSL and EX beams from the NF1611 on a scope (50-ohm input), the signal Vpp was ~200 mV.
  2. In the time that I was poking about, the level dropped to ~150mVpp. seemed suspicious.
  3. Thinking that this has to be related to the polarization mismatch between the interfering beams, I moved the input fibers (blue in Attachment #1) around, and saw the signal amplitude went up to 300mVpp, supporting my initial hypothesis.
  4. The question remains as to where the bulk of the polarization drift is happening. I had spent some effort making sure the input coupled beam to the fiber was well-aligned to one of the special axes of the fiber, and I don't think this will have changed since (i.e. the rotational orientation of the fiber axes relative to the input beam was fixed, since we are using the K6XS mounts with a locking screw for the input couplers). So I flexed the patch cables of the fiber beam splitters inside the BeatMouth, and saw the signal go as high as 700mVpp (the expected level given the values reported by the DC monitor).

This is a problem - such large shifts in the signal level means we have to leave sufficient headroom in the choice of RF amplifier gain to prevent saturation, whereas we want to boost the signal as much as possible. Moreover, this kind of operation of tweaking the fiber seating to increase the RF signal level is not repeatable/reliable. Options as I see it:

  1. Get a fiber BS that is capable of maintaining the beam polarization all the way through to the beat photodiode. I've asked AFW technologies (the company that made our existing fiber BS parts) if they supply such a device, and Andrew is looking into a similar component from Thorlabs.
    • These parts could be costly.
  2. Mix the beams in free space. We have the beam coming from EX to the PSL table, so once we mix the two beams, we can use either a fiber or free-space PD to read out the beatnote. 
    • This approach means we lose some of the advantages of the fiber based setup (e.g. frequent alignment of the free-space MM of the two interfering beams may be required).
    • Potentially increases sensitivity to jitter noise at the free-space/fiber coupling points
Quote:
    • An initial RF beat power level measurement yielded -5dBm, which is inconsistent with the DC monitor voltages, but I'm not sure what frequency the beat was at, will make a more careful measurement with a scope or the network analyzer.
Attachment 1: IMG_7384.JPG
IMG_7384.JPG
  14502   Fri Mar 29 21:00:06 2019 gautamUpdateALSBeatMouth with NF1611s installed
  • Newfocus 15V current limited supply was taken from bottom NE corner of the ITMY Oplev table to power the BeatMouth on the PSL table
  • BeatMouth was installed on top shelf on PSL table [Attachment #1].
  • Light levels in fibers were checked:
    • PSL: initially, only ~200uW / 4mW was coupled in. This was improved to 2.6mW/4mW (~65% MM) which was deemed sufficient for a first test), by tweaking the alignment of, and into the collimator.
    • EX: ~900uW measured at the PSL table. I remember the incident power being ~1mW. So this is pretty good.
  • Fibers hooked up to BeatMouth:
    • EX light only, DC mon of X PD reads -2.1V.
    • With PSL light, I get -4.6 V.
    • For these numbers, with the DC transimpedance of 10kohm and the RF transimpedance of 700 ohm, I expect a beat of ~0dBm
  • DC light level stability is being monitored by a temporarily hijacked PSL NPRO diagnostic Acromag channel. Main motivation is to confirm that the alignment to the special axes of the PM fibers is still good and we aren't seeing large tempreature-driven waveplate effects.
  • RF part of the circuit is terminated into 50ohms for now -
    • there is still a quesiton as to what is the correct RF amplifier to use in sending the signal to the 1Y3 rack.
    • An initial RF beat power level measurement yielded -5dBm, which is inconsistent with the DC monitor voltages, but I'm not sure what frequency the beat was at, will make a more careful measurement with a scope or the network analyzer.
    • We want the RF pre-amp to be:
      • Low noise, keeping this in mind
      • High enough gain to boost the V/Hz discriminant of the electronic delay line
      • Not too high gain that we run into compression / saturate some of the delay line electronics - specifically, the LO input of the LSC demod board has a Teledyne amp in the signal chain, and so we need to ensure the signal level there is <16dBm (nominal level is 10dBm).
      • I'm evaluating options...
  • At 1Y3:
    • I pulled out the delay-line enclosure, and removed the (superglued) resistive power splitters with the help of some acetone
    • The newly acquired power splitters (ZAPD-2-252-S+) were affixed to the front panel, in which I made some mounting holes.
    • The new look setup, re-installed at 1Y3, is shown in Attachment #2.
Attachment 1: IMG_7384.JPG
IMG_7384.JPG
Attachment 2: IMG_7385.JPG
IMG_7385.JPG
  14501   Fri Mar 29 15:47:58 2019 gautamUpdateAUXAUX laser fiber moved from AS table to PSL table

[anjali, gautam]

To facilitate the 1um MZ frequency stabilization project, I decided that the AUX laser was a better candidate than any of the other 3 active NPROs in the lab as (i) it is already coupled into a ~60m long fiber, (ii) the PSL table has the most room available to set up the readout optics for the delayed/non-delayed beams and (iii) this way I can keep working on the IR ALS system in parallel. So we moved the end of the fiber from the AS table to the SE corner of the PSL table. None of the optics mode-matching the AUX beam to the interferometer were touched, and we do not anticipate disturbing the input coupling into the fiber either, so it should be possible to recover the AUX beam injection into the IFO relatively easily.

Anjali is going to post detailed photos, beam layout, and her proposed layout/MM solutions later today. The plan is to use free space components for everything except the fiber delay line, as we have these available readily. It is not necessarily the most low-noise option, but for a first pass, maybe this is sufficient and we can start building up a noise budget and identify possible improvements.

The AUX laser remians in STANDBY mode for now. HEPA was turned up while working at the PSL table, and remains on high while Anjali works on the layout.

  14500   Fri Mar 29 11:43:15 2019 JonUpdateUpgradeFound c1susaux database bug

I found the current bias output channels, C1:SUS-<OPTIC>_<DOF>BiasAdj, were all pointed at C1:SUS-<OPTIC>_ULBiasSet for every degree of freedom. This same issue appeared in all eight database files (one per optic), so it looks like a copy-and-paste error. I fixed them to all reference the correct degree of freedom.

  14499   Thu Mar 28 23:29:00 2019 KojiUpdateSUSSuspension PD whitening and I/F boards modified for susaux replacement

Now the sus PD whitening bards are ready to move the back plane connectoresto the lower row and to plug the acromag interface board to the upper low.


Sus PD whitening boards on 1X5 rack (D000210-A1) had slow and fast channels mix in a single DIN96 connector. As we are going to use the rear-side backplane connector for Acromag access, we wanted to migrate the fast channel somewhere. For this purpose, the boards were modified to duplicate the fast signals to the lower DIN96 connector.

The modification was done on the back layer of the board (Attachment 1).
The 28A~32A and 28C~32C of P1 are connected to the corresponding pins of P2 (Attachment 2). The connections were thouroughly checked by a multimeter.

After the modification the boards were returned to the same place of the crate. The cables, which had been identified and noted before disconnection, were returned to the connectors.

The functionarity of the 40 (8sus*5ch) whitening switches were confimred using DTT one by one by looking at the transfer functions between SUS LSC EXC to the PD input filter IN1. All the switches showed the proper whitening in the measurments.

The PD slow mon (like C1:SUS-XXX_xxPDMon) channels were also checked and they returned to the values before the modification, except for the BS UL PD. As the fast version of the signal returned to the previous value, the monitor circuit was suspicious. Therefore the opamp of the monitor channels (LT1125) were replaced and the value came back to the previous value (attachment 3).

 

Attachment 1: IMG_7474.JPG
IMG_7474.JPG
Attachment 2: D000210_backplane.pdf
D000210_backplane.pdf
Attachment 3: Screenshot_from_2019-03-28_23-28-23.png
Screenshot_from_2019-03-28_23-28-23.png
  14498   Thu Mar 28 19:40:02 2019 gautamUpdateALSBeatMouth with NF1611s assembled

Summary:

The parts I was waiting for arrived. I finished the beat mouth assembly, and did some characterization. Everything looks to be working as expected.

Details:

Attachment #1: Photo of the front panel. I am short of two fiber mating sleeves that are compatible with PM fibers, but those are just for monitoring, so not critical to the assembly at this stage. I'll ask Chub to procure these.

Attachment #2: Photo of the inside of the BeatMouth. I opted to use the flexible RG-316 cables for all the RF interconnects. Rana said these aren't the best option, remains to be seen if interference between cables is an issue. If so, we can replace them with RG-58. I took the opportunity to give each fiber beam splitter its own spool, and cleaned all the fiber tips.

Attachment #3: Transfer function measurement. The PDFR setup behind 1X5/1X6 was used. I set the DC current to the laser to 30.0 mA (as read off the display of the current source), which produced ~400uW of light at the fiber coupled output of the diode laser. This was injected into the "PSL" input coupler of the BeatMouth, and so gets divided down to ~100 uW by the time it reaches the PDs. From the DC monitor values (~430mV), the light hitting the PDs is actually more consistent with 60uW, which is in agreement with the insertion loss of the fiber beamsplitters, and the mating sleeves.

The two responses seem reasonably well balanced (to within 20% - do we expect this to be better?). Even though judging by the DC monitor, there was more light incident on the Y PD than on the X PD, the X response was actually stronger than the Y. 

I also took the chance to do some other tests:

  • Inject light into the "X(Y)-ARM" input coupler of the Beat Mouth - confirmed that only the X(Y) NF1611's DC monitor output showed any change. The DC light level was ~1V in this condition, which again is consistent with expected insertion losses as compared to the "PSL" input case, there is 1 less fiber beamsplitter and mating sleeve.
  • Injected light into each of the input couplers, looked at the interior of the BeatMouth with an IR viewer for evidence of fiber damage, and saw none. Note that we are not doing anything special to dump the light at the unused leg of the fiber beamsplitter (which will eventually be a monitor port). Perhaps, nominally, this port should be dumped in some appropriate way.

Attachment #4: Dark Noise analysis. I used a ZHL-500-HLN+ to boost the PD's dark noise above the AG4395's measurement noise floor. The measured noise level seems to suggest either (i) the input-referred current noise of the PD circuitry is a little lower than the spec of 16 pA/rtHz (more like 13 pA/rtHz) or (ii) the transimpedance is lower than the spec of 700 V/A (more like 600 V/A). Probably some combination of the two. Seems reasonable to me.

Next steps:

The optical part of the ALS detection setup is now complete. The next step is to measure the ALS noise with this sysytem. I will use the X arm for this purpose (I'd like to make the minor change of switching the existing resistive power splitter at the delay line to the newly acquired splitters which have 3dB lower insertion loss). 

Attachment 1: IMG_7381.JPG
IMG_7381.JPG
Attachment 2: IMG_7382.JPG
IMG_7382.JPG
Attachment 3: relTF_schem.pdf
relTF_schem.pdf
Attachment 4: darkNoise.pdf
darkNoise.pdf
  14497   Tue Mar 26 18:35:06 2019 JonUpdateUpgradeModbus IOC is running on c1susaux2

Thanks to new info from Johannes, I was able to finish setting up the modbus IOC on c1susaux2. It turns out the 17 Acromags draw ~1.9 A, which is way more than I had expected. Hence the reason I had suspected a short. Adding a second DC supply in parallel solves the problem. There is no issue with the wiring.

With the Acromags powered on, I carried out the following:

  • Confirmed c1susaux2 can communicate with each Acromag at its assigned IP address
  • Modified the EPICS .cmd file to point to the local modbus installation (not the remote executable on /cvs/cds)
  • Debugged several IOC initialization errors. All were caused by minor typos in the database files.
  • Scripted the modbus IOC to launch as a systemd service (will add implementation details to the documentation page)

The modbusIOC is now running as a peristent system service, which is automatically launched on boot and relaunched after a crash. I'm able to access a random selection of channels using caget.

What's left now is to finish the Acromag-to-feedthrough wiring, then test/calibrate each channel.

  14496   Tue Mar 26 04:25:13 2019 JohannesUpdateUpgradec1susaux upgrade plan
Quote:

Main TODO items

  • Debug issue with Acromag DC power wiring
  • Complete wiring from chassis feedthroughs to Acromag terminals, following this wiring diagram
  • Check/set the configuration of each Acromag unit using the software on the Windows laptop
  • Set the analog channel calibrations in the EPICS database file
  • Test each channel ex situ. Chub and I discussed an idea to use two DB-37F breakout boards, with the wiring between the board terminals manually set. One DAC channel would be calibrated and driven to test other ADC channels. A similar approach could be used for the digital input/output channels.

Just a few remarks, since I heard from Gautam that c1susaux is next in line for upgrade.

All units have already been configured with IP addresses and settings following the scheme explained on the slow controls wiki page. I did this while powering the units in the chassis, so I'm not sure where the short is coming from. Is the power supply maybe not sourcing enough current? Powering all units at the same time takes significant current, something like >1.5 Amps if I remember correctly. These are the IPs I assigned before I left:

Acromag Unit IP Address
C1SUSAUX_ADC00 192.168.115.20
C1SUSAUX_ADC01 192.168.115.21
C1SUSAUX_ADC02 192.168.115.22
C1SUSAUX_ADC03 192.168.115.23
C1SUSAUX_ADC04 192.168.115.24
C1SUSAUX_ADC05 192.168.115.25
C1SUSAUX_ADC06 192.168.115.26
C1SUSAUX_ADC07 192.168.115.27
C1SUSAUX_ADC08 192.168.115.28
C1SUSAUX_ADC09 192.168.115.29
C1SUSAUX_DAC00 192.168.115.40
C1SUSAUX_DAC01 192.168.115.41
C1SUSAUX_DAC02 192.168.115.42
C1SUSAUX_DAC03 192.168.115.43
C1SUSAUX_BIO00 192.168.115.60
C1SUSAUX_BIO01 192.168.115.61
C1SUSAUX_BIO02 192.168.115.62

I used black/white twisted-pair wires for A/D, red/white for D/A, and green/white for BIO channels. I found it easiest to remove the blue terminal blocks from the Acromag units for doing the majority of the wiring, but wasn't able to finish it. I had also done the analog channel calibrations using the windows untility using multimeters and one of the precision voltage sources I had brought over from the Bridge labs, but it's probably a good idea to check it and correct if necessary. I also recommend to check that the existing wiring particularly for MC1 and MC2 is correct, as I had swapped their order in the channel assignment in the past.

While looking through the database files I noticed two glaring mistakes which I fixed:

  1. The definition of C1SUSAUX_BIO2 was missing in /cvs/cds/caltech/target/c1susaux2/C1SUSAUX.cmd. I added it after the assignments for C1SUSAUX_BIO1
  2. Due to copy/paste the database files /cvs/cds/caltech/target/c1susaux2/C1_SUS-AUX_<OPTIC>.db files were still pointing to C1AUXEX. I overwrote all instances of this in all database files with C1SUSAUX.

 

  14495   Mon Mar 25 10:21:05 2019 JonUpdateUpgradec1susaux upgrade plan

Now that the Acromag upgrade of c1vac is complete, the next system to be upgraded will be c1susaux. We chose c1susaux because it is one of the highest-priority systems awaiting upgrade, and because Johannes has already partially assembled its Acromag replacement (see photos below). I've assessed the partially-assembled Acromag chassis and the mostly-set-up host computer and propose we do the following to complete the system.

Documentation

As I go, I'm writing step-by-step documentation here so that others can follow this procedure for future systems. The goal is to create a standard procedure that can be followed for all the remaining upgrades.

Acromag Chassis Status

The bulk of the remaining work is the wiring and testing of the rackmount chassis housing the Acromag units. This system consists of 17 units: 10 ADCs, 4 DACs, and 3 digitial I/O modules. Johannes has already created a full list of channel wiring assignments. He has installed DB37-to-breakout board feedthroughs for all the signal cable connections. It looks like about 40% of the wiring from the breakout boards to Acromag terminals is already done.

The Acromag units have to be initially configured using the Windows laptop connected by USB. Last week I wasn't immediately able to check their configuration because I couldn't power on the units. Although the DC power wiring is complete, when I connected a 24V power supply to the chassis connector and flipped on the switch, the voltage dropped to ~10V irrespective of adjusting the current limit. The 24V indicator lights on the chassis front and back illuminated dimly, but the Acromag lights did not turn on. I suspect there is a short to ground somewhere, but I didn't have time to investigate further. I'll check again this week unless someone else looks at it first.

Host Computer Status

The host computer has already been mostly configured by Johannes. So far I've only set up IP forwarding rules between the martian-facing and Acromag-facing ethernet interfaces (the Acromags are on a subnet inaccessible from the outside). This is documented in the link above. I also plan to set up local installations of modbus and EPICS, as explained below. The new EPICS command file (launches the IOC) and database files (define the channels) have already been created by Johannes. I think all that remains is to set up the IOC as a persistent system service.

Host computer OS

Recommendation from Keith Thorne:

For CDS lab-wide, Jamie Rollins and Ryan Blair have been maintaining Debian 8 and 9 repos with some of these.  
They have somewhat older EPICS versions and may not include all the modules we have for SL7.
One worry is whether they will keep up Debian 9 maintained, as Debian 10 is already out.

I would likely choose Debian 9 instead of Ubuntu 18.04.02, as not sure of Ubuntu repos for EPICS libraries.

Based on this, I propose we use Debian 9 for our Acromag systems. I don't see a strong reason to switch to SL7, especially since c1vac and c1susaux are already set-up using Debian 8. Although Debian 8 is one version out of date, I think it's better to get a well-documented and tested procedure in place before we upgrade the working c1vac and c1susaux computers. When we start building the next system, let's install Debian 9 (or 10, if it's available), get it working with EPICS/modbus, then loop back to c1vac and c1susaux for the OS upgrade.

Local vs. central modbus/EPICS installation

The current convention is for all machines to share a common installation which is hosted on the /cvs/cds network drive. This seems appealing because only a single central EPICS distribution needs to be maintained. However, from experience attempting this on c1vac, I'm convinced this is a bad design for the new Acromag systems.

The problem is that any network outage, even routine maintenance or brief glitches, wreaks havoc on Acromags set up this way. When the network is interrupted, the modbus executable disappears mid-execution, crashing the process and hanging the OS (I think related to the deadlocked NFS mount), so that the only way to recover is to manually power-cycle. Still worse, this can happen silently (channel values freeze), meaning that, e.g., watchdog protections might fail.

To avoid this, I'm planning to install a local EPICS distribution from source on c1susaux, just as I did for c1vac. This only takes a few minutes to do, and I will include the steps in the documented procedure. Building from source also better protects against OS-dependent buginess.

Main TODO items

  • Debug issue with Acromag DC power wiring
  • Complete wiring from chassis feedthroughs to Acromag terminals, following this wiring diagram
  • Check/set the configuration of each Acromag unit using the software on the Windows laptop
  • Set the analog channel calibrations in the EPICS database file
  • Test each channel ex situ. Chub and I discussed an idea to use two DB-37F breakout boards, with the wiring between the board terminals manually set. One DAC channel would be calibrated and driven to test other ADC channels. A similar approach could be used for the digital input/output channels.
Attachment 1: IMG_3136.jpg
IMG_3136.jpg
Attachment 2: IMG_3138.jpg
IMG_3138.jpg
Attachment 3: IMG_3137.jpg
IMG_3137.jpg
  14494   Thu Mar 21 21:50:31 2019 ranaUpdateVACProtection against AC power loss

agreed - we need all pumps on UPS for their safety and also so that we can spin them down safely. Can you and Chub please find a suitable UPS?

Quote:

However, I discovered that TP1---the pump that might be most damaged by a sudden power failure---is not on the UPS. It's plugged directly into a 240V outlet along the wall. This is because the current UPS doesn't have any 240V sockets. I'd recommend we get one that can handle all the turbo pumps.

  14493   Thu Mar 21 18:36:59 2019 JonOmnistructureUpgradeVacuum Controls Switchover Completed

Updated vac channel list is attached. There are several new ADC channels.

Quote:

Hardware & Channel Assignments

All of the new hardware is now permanently installed in the vacuum rack. This includes the SuperMicro rack server (c1vac), the IOLAN serial device server, a vacuum subnet switch, and the Acromag chassis. Every valve/pump signal cable that formerly connected to the VME bus through terminal blocks has been refitted with a D-sub connector and screwed directly onto feedthroughs on the Acromag chassis.

The attached pdf contains the master list of assigned Acromag channels and their wiring.

Attachment 1: 40m_Vacuum_Acromag_Channels_20190321.pdf
40m_Vacuum_Acromag_Channels_20190321.pdf 40m_Vacuum_Acromag_Channels_20190321.pdf 40m_Vacuum_Acromag_Channels_20190321.pdf
  14492   Thu Mar 21 18:09:36 2019 KojiUpdateCDSdb file preparation for acromag c1susaux

I have updated the google doc spreadsheet to indicate the required action for the new dbfile generation.

There are three types of actions:

1. COPY - Just duplicate the old EPICS db entry. This is for soft channels, calc channels.
2. DELETE - Delete the entry for some physical channels that will not be implemented on Acromag (oplev, dewhitening mon, AI monitor, etc)
3. REPLACE - For the physical channels, we want to replace the port names.

The blue part of the spreadsheet indicates the action for each channel. If it is a physical channel, the assigned module and the channel are indicated there. What we still want to do is to use the these information for generating the port name which looks like "@asynMask(C1VAC_XT1221A_ADC 1 -16)MODBUS_DATA".

The links to the spreadsheets can be found on 40m wiki: https://wiki-40m.ligo.caltech.edu/CDS/SlowControls/c1susaux

Attachment 1: Screen_Shot_2019-03-21_at_18.06.53.png
Screen_Shot_2019-03-21_at_18.06.53.png
  14491   Thu Mar 21 17:22:52 2019 JonUpdateVACMore vac controls upgrades

I've converted all the vac control system code to run on Python 3.4, the latest version available through the Debian package manager. Note that these codes now REQUIRE Python 3.x. We decided there was no need to preserve Python 2.x compatibility. I'm leaving the vac system returned to its nominal state ("vacuum normal + RGA").

Quote:

The vac controls system is going down for migration from Python 2.7 to 3.4. Will advise when it is back up.

 

  14490   Thu Mar 21 12:46:22 2019 JonUpdateVACMore vac controls upgrades

The vac controls system is going down for migration from Python 2.7 to 3.4. Will advise when it is back up.

  14489   Wed Mar 20 20:07:22 2019 JonUpdateVACDoing vac controls work

Work is completed and the vac system is back in its nominal state.

Quote:

I'm rebooting the IOLAN server to load new serial ports. The interlocks might trip when the pressure gauge readbacks cut out.

 

  14488   Wed Mar 20 19:26:25 2019 JonUpdateVACProtection against AC power loss

Today I implemented protection of the vac system against extended power losses. Previously, the vac controls system (both old and new) could not communicate with the APC Smart-UPS 2200 providing backup power. This was not an issue for short glitches, but for extended outages the system had no way of knowing it was running on dwindling reserve power. An intelligent system should sense the outage and put the IFO into a controlled shutdown, before the batteries are fully drained.

What enabled this was a workaround Gautam and I found for communicating with the UPS serially. Although the UPS has a serial port, neither the connector pinout nor the low-level command protocol are released by APC. The only official way to communicate with the UPS is through their high-level PowerChute software. However, we did find "unofficial" documentation of APC's protocol. Using this information, I was able to interface the the UPS to the IOLAN serial device server. This allowed the UPS status to be queried using the same Python/TCP sockets model as all the other serial devices (gauges, pumps, etc.). I created a new service called "serial_UPS.service" to persistently run this Python process like the others. I added a new EPICS channel "C1:Vac-UPS_status" which is updated by this process.

With all this in place, I added new logic to the interlock.py code which closes all valves and stops all pumps in the event of a power failure. To be conservative, this interlock is also tripped when the communications link with the UPS is disconnected (i.e., when the power state becomes unknown). I tested the new conditions against both communication failure (by disconnecting the serial cable) and power failure (by pressing the "Test" button on the UPS front panel). This protects TP2 and TP3. However, I discovered that TP1---the pump that might be most damaged by a sudden power failure---is not on the UPS. It's plugged directly into a 240V outlet along the wall. This is because the current UPS doesn't have any 240V sockets. I'd recommend we get one that can handle all the turbo pumps.

For future reference:

Pin 1: RxD

Pin 2: TxD

Pin 5: GND

Standard: RS-232

Baud rate: 2400

Data bits: 8

Parity: none

Stop bits: 1

Handshaking: none

 

 

Attachment 1: IMG_3146.jpg
IMG_3146.jpg
  14487   Wed Mar 20 12:31:30 2019 JonUpdateVACDoing vac controls work

I'm rebooting the IOLAN server to load new serial ports. The interlocks might trip when the pressure gauge readbacks cut out.

  14486   Mon Mar 18 20:22:28 2019 gautamUpdateALSALS stability test

I'm running a test to see how stable the EX green lock is. For this purpose, I've left the slow temperature tuning servo on (there is a 100 count limiter enabled, so nothing crazy should happen).

  14485   Mon Mar 18 18:10:14 2019 KojiSummaryGeneralTask items and priority

[Gautam/Chub/Koji] ~ Mini discussion

Maintenance / Upgrade Items

(Priority high to low)

  • TT/IO suspension upgrade (solidworks work) -> order components -> TT characterization
  • Acromag upgrade c1susaux
    • Produce spread sheetfor DB files. Learn new format of the DB file with Acromag. Develop a python code for the DB file generation (Jon->Koji)
  • Satellite Box upgrade
    • Rack mount? Front panel DB connectors. New circuits (PD-LED)
       
  • Acromag iscaux1/2 & isc whitening upgrade
     
  • new RC mirror characterization -> installation
  14484   Mon Mar 18 17:06:12 2019 gautamUpdateOptical LeversITMY HeNe replaced

Oplev HeNe was replaced this afternoon. We did some HeNe shuffling:

  1. A new HeNe was being used for the fiber illumination demo at EX. We took that out and decided to use it as the new ITMX HeNe. It had 2.6mW output at 632nm (measured with the Ophir power meter)
  2. Old ETMY HeNe was used for fiber illumination demo.
  3. Old ITMX HeNe was putting out no light - it will be disposed.

Attachment #1 shows the RIN and Attachment #2 and #3 show the PIT and YAW TFs with the new HeNe.

The ITMX Oplev path is still not great - the ingoing beam is within 2mm of clipping on a 2" lens used in the POX path, and there is a bunch of scattered red light everywhere. We should take the opportunity when the chamber is open to try and have a better layout (it may be tricky to optize without touching the two in-vacuum steering optics).

Quote:

I'll ask Chub to replace it this afternoon.

Attachment 1: OLRIN.pdf
OLRIN.pdf
Attachment 2: OL_PIT.pdf
OL_PIT.pdf
Attachment 3: OL_YAW.pdf
OL_YAW.pdf
  14483   Mon Mar 18 12:27:42 2019 gautamUpdateGeneralIFO status
  1. c1iscaux2 VME crate is damaged - see Attachment #1. 
    • It is not generating the 12V supply voltage, and so nothing in the crate works.
    • Tried resetting via front panel button, power cycling by removing power cable on rear, all to no effect.
    • Tried pulling out all cards and checking if there was an internal short that was causing the failure - looks like the problem is with the crate itself.
    • Not sure how long this machine has been unresponsive as we don't have any readback of the status of the eurocrate machines.
    • Not a showstopper, mainly we can't control the whitening settings for AS55, REFL55, REFL165 and ALSY. 
    • Acromag installation schedule should be accelerated.
    • * Koji reminded me that \text{VME crate} \ \neq \ \text{eurocrate}. The former is what is used for the slow machines, the latter is what is used for holding the iLIGO style electronics boards.
  2. ITMX oplev is dead - see Attachment #2.
    • Lasted ~3 years (installed March 2016).
    • I confirmed that no light is coming out of the laser head on the optical table.
    • I'll ask Chub to replace it this afternoon.
  3. c1susaux is unresponsive
    • I didn't reboot it as I didn't want to spend some hours freeing ITMY. 
    • At some point we will have to bite the bullet and do it.
  4. Input pointing is still not stable
    • I aligned the input pointing using TT1/TT2 to maximize TRX/TRY before lunch, but in 1 hour, the pointing has already drifted.
  5. POX/POY locking is working okay. TRX has large low-frequency fluctuations because of ITMX not having an Oplev servo, should be rectified once we swap out the HeNe.

The goal for this week is to test out the ALS system, so this is kind of a workable state since POX/POY locking is working. But the number of broken things is accumulating fast.

Attachment 1: IMG_7343.JPG
IMG_7343.JPG
Attachment 2: ITMXOL.png
ITMXOL.png
  14482   Sun Mar 17 21:06:17 2019 AnjaliUpdateALSAmplifier characterisation

The goal was to characterise the new amplifier (AP1053). For a practice, I did the characterisation of the old amplifier.This test is similar to that reported in Elog ID 13602.

  • Attachment #1 shows the schematic of the setup for gain characterisation and Attachment #2 shows the results of gain characterisation. 
  • The gain measurement is comparable with the previous results. From the data sheet, 10 dB gain is guaranteed in the frequency range 10-450 MHz. From our observation, the gain is not flat pver this region. We have measured a maximum gain of 10.7 dB at 6 MHz and it has then decreased upto 8.5 dB at 500 MHz
  • Attachement #3 shows the schematic of the setup for the noise characterisation and Attachment # 4 shows the results of noise measurment. 
  • The noise measurement doesn't look fine. We probably have to repeat this measurement.
Attachment 1: Gain_measurement.pdf
Gain_measurement.pdf
Attachment 2: Amplifier_gain.pdf
Amplifier_gain.pdf
Attachment 3: noise_measurement.pdf
noise_measurement.pdf
Attachment 4: noise_characterisation.pdf
noise_characterisation.pdf
  14481   Sun Mar 17 13:35:39 2019 AnjaliUpdateALSPower splitter characterization

We characterized the power splitter ( Minicircuit- ZAPD-2-252-S+). The schematic of the measurement setup is shown in attachment #1. The network/spectrum/impedance analyzer (Agilent 4395A) was used in the network analyzer mode for the characterisation. The RF output is enabled in the network analyser mode. We used an other spliiter (Power splitter #1) to splitt the RF power such that one part goes to the network analzer and the other part goes to the power spliiter (Power splitter #2) . We are characterising power splitter #2 in this test. The characterisation results and comparison with the data sheet values are shown in Attachment # 2-4.

Attachment #2 : Comparison of total loss in port 1 and 2

Attachment #3 : Comparison of amplitude unbalance

Attachment #4 : Comparison of phase unbalance

  • From the data sheet: the splitter is wideband, 5 to 2500 MHz, useable from 0.5 to 3000 MHz. We performd the measurement from 1 MHz to 500 MHz (limited by the band width of the network analyzer).
  • It can be seen from attachment #2 and #4 that there is a sudden increase below ~11 MHz. The reason for this is not clear to me
  • The mesured total loss value for port 1 and port 2 are slightly higher than that specified in the data sheet.From the data sheet, the maximum loss in port 1 and port 2 in the range at 450 MHz are 3.51 dB and 3.49 dB respectively. The measured values are 3.61 dB and 3.59 dB respectively for port 1 and port 2, which is higher than the values mentioed in the data sheet. It can also be seen from attachment #1 (b) that the expected trend in total loss with frequency is that the loss is decreasing with increase in frequency and we are observing the opposite trend in the frequency range 11-500 MHz. 
  • From the data sheet, the maximum amplitude balance in the 5 MHz-500 MHz range is 0.02 dB and the measured maximum value is 0.03 dB
  • Similary for the phase unbalance, the maximum value specified by the data sheet in the 5 MHz- 500 MHz range is 0.12 degree and the measurement shows a phase unbalance upto 0.7 degree in this frequency range
  • So the observations shows that the measured values are slighty higher than that specified in the data sheet values.
Attachment 1: Measurement_setup.pdf
Measurement_setup.pdf
Attachment 2: Total_loss.pdf
Total_loss.pdf
Attachment 3: Amplitude_unbalance.pdf
Amplitude_unbalance.pdf
Attachment 4: Phase_unbalance.pdf
Phase_unbalance.pdf
  14480   Sun Mar 17 00:42:20 2019 gautamUpdateALSNF1611 cannot be shot-noise limited?

Summary:

Per the manual (pg12) of the NF 1611 photodiode, the "Input Noise Current" is 16 pA/rtHz. It also specifies that for "Linear Operation", the max input power is 1 mW, which at 1um corresponds to a current shot noise of ~14 pA/rtHz. Therefore,

  1. This photodiode cannot be shot-noise limited if we also want to stay in the spec-ed linear regime.
  2. We don't need to worry so much about the noise figure of the RF amplifier that follows the photodiode. In fact, I think we can use a higher gain RF amplifier with a slightly worse noise figure (e.g. ZHL-3A) as we will benefit from having a larger frequency discriminant with more RF power reaching the delay line.

Details:

Attachment #1: Here, I plot the expected voltage noise due to shot noise of the incident light, assuming 0.75 A/W for InGaAs and 700V/A transimpedance gain. 

  • For convenience, I've calibrated on the twin axes the current shot noise (X) and equivalent amplifier noise figure at a given voltage noise, assuming a 50 ohm system (Y).
  • The 16 pA/rtHz input current noise exceeds the shot noise contribution for powers as high as 1 mW.
  • Even at 0.5 mW power on the PD, we can use the ZHL-3A rather than the Teledyne:
    • This calculation was motivated by some suspicious features in the Teledyne amplifier gain, I will write a separate elog about that. 
    • For the light levels we have, I expect ~3dBm RF signal from the photodiode. With the 24dB of gain from the ZHL-3A, the signal becomes 27dBm, which is smaller (but close to) the spec-ed max output of the ZHL-3A, which is 29.5 dBm. Is this too close to the edge?
    • I will measure the gain/noise of the ZHL-3A to get a better answer to these questions.
  • If in the future we get a better photodiode setup that reaches sub-1nV/rtHz (dark/electronics) voltage noise, we may have to re-evaluate what is an appropriate RF amplifier.
Attachment 1: PDnoise.pdf
PDnoise.pdf
ELOG V3.1.3-